OpenStudy (anonymous):

solve. 3/x+4=1/x-1 can someone please show me how to solve this in all the steps.

OpenStudy (solomonzelman):

\[\frac{3}{x+4}=\frac{1}{x-1}\]you can therefore say that,\[\frac{x+4}{3}=\frac{x-1}{1}\]\[\frac{x+4}{3}=x-1\]\[\frac{x+4}{3}\color{red} {\times 3}=(x-1)\color{red} {\times 3}\]

OpenStudy (solomonzelman):

\[x+4=3x-3\]\[x+4\color{red} {-x+3}=3x-3\color{red} {-x+3}\]\[7=2x\]\[x=3.5~~~~~or~~~~~3\frac{1}{2}\]

OpenStudy (anonymous):

thank you so much for your help but can you also write that as x=7/2?

OpenStudy (anonymous):

x=7/2?

OpenStudy (solomonzelman):

Yes, and anytime!